### v1

# DiPT: Enhancing LLM Reasoning through Diversified Perspective-Taking

Abstract
Existing work on improving language model reasoning typically explores a single solution path, which can be prone to errors. Inspired by perspective-taking in social studies, this paper introduces DiPT, a novel approach that complements current reasoning methods by explicitly incorporating diversified viewpoints. This approach allows the model to gain a deeper understanding of the problem’s context and identify the most effective solution path during the inference stage. Additionally, it provides a general data-centric AI recipe for augmenting existing data to improve their quality for fine-tuning. Our empirical results demonstrate that DiPT can be flexibly integrated into existing methods that focus on a single reasoning approach, enhancing their reasoning performance and stability when presented with paraphrased problems. Furthermore, we illustrate improved context understanding by maintaining the model’s safe outputs against "jailbreaking" prompts intentionally designed to bypass safeguards built into deployed models. Lastly, we show that fine-tuning with data enriched with diverse perspectives can boost the reasoning capabilities of the model compared to fine-tuning with raw data alone.
Keywords Reasoning  $\cdot$ Perspective-Taking  $\cdot$ LLM
1 Introduction
Correct reasoning steps are important for language models to achieve high performance on many tasks, such as commonsense reasoning, question answering, and mathematical problem-solving (Wei et al., 2022; Kojima et al., 2022; Suzgun et al., 2022). One way to elicit reasoning is through the chain-of-thought (CoT) method Wei et al. (2022); Kojima et al. (2022), which asks the model to provide step-by-step reasoning. Another approach encourages the model to provide similar problems Yasunaga et al. (2024) as the query, indirectly compelling the model to first understand the original query. Similarly, repeating and rephrasing the query Deng et al. (2023); Mekala et al. (2023) requires the model to first understand the problem and then modify the query into its own words. This rephrasing might help simplify the problem for the model. Additionally, reasoning can be generated by indirectly providing reasoning examples in demonstrations, referred to as in-context learning (ICL) Brown et al. (2020); Min et al. (2022); Xie et al. (2021).
While these methods have demonstrated significant performance improvements, language models are still prone to errors due to incorrect context understanding or analytical steps. Furthermore, they are subject to instability when requests are paraphrased. This instability is particularly concerning in the context of adversarial prompts, where recent research (Zou et al., 2023; Zeng et al., 2024) has shown that adversaries can intentionally rewrite prompts to coax safety-aligned language models into generating objectionable content that they would not generate otherwise. Although the exact source of these errors is a subject of active research Kalai and Vempala (2024), we observe a commonality among these methods: they often generate an answer to the problem by considering only a single solution path, or perspective, by default. Figure 3 illustrates an example of an arithmetic question that is consistently answered incorrectly even by the most capable models (such as ChatGPT, Gemini as of date June 15, 2024). In this example, the direct application of existing methods, such as chain-of-thought, adopts a uniform strategy to answer it, leading to the wrong answer.
Figure 1: An illustration of enhancing current reasoning methods with perspective-taking.
On the other hand, in social studies, diversified perspective-taking—referring to the process where individuals deliberately consider multiple viewpoints when analyzing a problem—has demonstrated effectiveness in enhancing problem-solving performance Wang et al. (2006); LaRusso et al. (2016) and mitigating erroneous perceptions caused by a single biased perspective Galinsky and Moskowitz (2000); McCrudden et al. (2017). Inspired by this success, we hypothesize applying this technique to existing reasoning methods can enhance language models’ reasoning capabilities.
Technical Contributions. To this end, we introduce DiPT, a framework that extends reasoning methods with Diversified Perspective-Taking. This framework can be applied to both the inference and training stages. In the inference stage, it explicitly instructs the model to analyze alternative perspectives on the given problem, such as considering different options in multiple-choice questions or evaluating various solution methods for open-ended questions. In the training stage, it serves as a general recipe for improving data quality for fine-tuning, aligning with the principles of data-centric AI. Specifically, it uses an off-the-shelf model prompted to provide rationales from different perspectives, thereby enriching the information within the existing instruction-tuning datasets. Intuitively, fine-tuning on such enriched datasets encourages the model to go beyond memorizing input-output associations to understanding the skills and knowledge relevant to answering questions.
Empirical Takeaways. We performed extensive experiments to understand the potential unlocked by diversified perspective-taking in LLMs. The key findings are summarized as follows:
  • DiPT can be flexibly integrated into existing inference-time reasoning-enhancement methods, consistently improving accuracy by up to 6% and reducing inconsistency caused by questions’ paraphrases. Notably, it encourages self-correction, allowing the model to rectify errors made at a single solution path by corroborating answers from alternative perspectives.
  • DiPT leads to improved context understanding. We demonstrate this by applying DiPT as a moderation method to protect the system from jailbreaking queries that could elicit harmful content while maintaining utility on general queries.
  • DiPT leads to improved data quality for fine-tuning. A case study on fine-tuning revealed that using chain-of-thought data enriched with perspective-taking consistently yields improvements across various models and domains, compared to fine-tuning on raw data or data augmented with single-perspective chain-of-thought explanations. These improvements were observed both when evaluating on the same distribution data as the training set and when applied to different datasets within the same domain.
  • Our framework enables to effectively detect potential errors in datasets. We observe a wide range of labeling errors in commonly used datasets in the current literature. This finding highlights the need for high-quality datasets to improve the interpretation of results and the reliability of benchmarks.
2 Related Work
Improving Reasoning in the Inference Time.
Numerous single-prompt (0-shot) methods have emerged to improve the model’s reasoning capabilities. One such method is (automatic/0-shot) chain-of-thought (CoT) Wei et al. (2022); Kojima et al. (2022), which instructs the model to provide a step-by-step explanation of the answer. This can be achieved by either incorporating examples with such explanations or by introducing an additional sentence in the prompt, "Let’s think Step by Step." Plan-and-solve (PS) method Wang et al. (2023a) is the extension of the CoT reasoning, which asks the model to first come up with the plan before solving the problem in a step-by-step manner. Recent work also derives theoretical analysis Feng et al. (2023) explaining how the transformers with chain-of-thought reasoning can solve mathematical problems that otherwise would not be possible without outputting the reasoning by the model. Another line of work Mekala et al. (2023); Deng et al. (2023) attempts to involve the model to simplify the query before actually solving the problem by asking the model to rephrase the query in the model’s simplified language. With a simplified query, the model can better understand the problem and proceed to solve the task. Analogical reasoners Yasunaga et al. (2024), on the other hand, instructs the LLM to self-generate similar examples to the query as demonstrations and then solve the problem. Overall, the common limitation of these methods is that they do not regulate how reasoning should be performed and, by default, adopt a single solution path. This can be attributed to various factors, such as the simplicity and computational efficiency of generating a single solution path, the lack of explicit rewards for diversity in the reasoning process in current evaluation metrics, and the assumption that a correct solution path indicates sufficient problem understanding. While investigating the mechanisms that encourage the generation of a single solution path is beyond the scope of this paper, we focus on studying the empirical benefits of incorporating multiple solution paths for both inference and training stages. Improving the effectiveness of single prompting naturally involves incorporating multiple prompts, such as CoT self-consistency Wang et al. (2022), least-to-most prompting Zhou et al. (2022), (probabilistic) tree-of-thoughts (ToT) Yao et al. (2023); Cao et al. (2023), or graph-of-thoughts Besta et al. (2023). These methods enhance responses by leveraging diverse model outputs. While diversified perspective-taking shows promise in improving reasoning based on multiple prompts by increasing the accuracy of individual prompts, this paper focuses on integrating diverse perspectives into zero-shot methods as a proof of concept.
Improving Data Quality for Targeted Instruction Tuning.
Recent advancements in instruction tuning have enhanced the task-specific capabilities of large language models (LLMs) (Peng et al., 2023; Zhang et al., 2023). Existing work has developed various techniques to identify the most relevant data from these extensive datasets to effectively develop specific capabilities (Albalak et al., 2023; Xia et al., 2023, 2024; Xie et al., 2024; Kang et al., 2024). However, these methods all focus on pruning samples to distill the most informative pieces from a dataset. Instead, we explore how to enrich the information content of each sample and examine its impact. Others investigate rewriting individual samples to improve their quality, such as incorporating in-context examples (Longpre et al., 2023) and chain-of-thought reasoning into the instruction tuning dataset (Kim et al., 2023; Chai et al., 2024). By contrast, we explore whether incorporating perspective-taking data can further enhance instruction tuning performance.
3 DiPT: Diversified Perspective-taking
Now, we delve into the specifics of incorporating diversified perspective-taking into the inference and fine-tuning stages of language models.
3.1 DiPT as an Inference-Time Reasoning Enhancement Tool
The key idea behind DiPT applied to inference time is to prompt the model to consider multiple perspectives or solution paths for a given problem before attempting to solve it. This explicit consideration of diverse perspectives is a crucial distinction between DiPT and previous approaches where perspective-taking is not explicitly called for and does not occur most of the time in the generation.
In a standard case, for a given problem, a perspective would be implicitly invoked and a reasoning method would be applied to solve it (e.g., with CoT reasoning). In the case of perspective-taking, multiple perspectives would be proposed, and we would apply the reasoning method to each of the given perspectives as seen in Figure 1. To instantiate perspectives in different problem contexts, we propose the following strategies.
For multiple-choice problems, DiPT prompts the model to re-think each option as a potential solution. In doing so, the model is encouraged to thoroughly consider each choice, rather than simply selecting the first most likely option based on next-word prediction, which might be subject to spurious correlations existing in the training data. For instance, consider the following text: “The backyard battles you staged with your green plastic army men were more exciting and almost certainly made more sense.” GPT-4-0613 labeled this text as having a positive sentiment, potentially due to the presence of certain words such as “exciting” and “make more sense,” which often appear in positive reviews. However, by explicitly prompting the model to think from the perspective that this review might be negative, it can better grasp the text’s negative tone and correct its answer. This example demonstrates how DiPT can help the model overcome misleading cues and better understand the overall context to make more accurate predictions. We refer the reader to Appendix A.5 and Appendix B.2 for detailed prompts and response.
For open-ended questions or free-text generation problems, DiPT prompts the model to consider different methods or approaches to solve the problem. For example, the question “What is the sum of all numbers between $-27\leq x<27$?” is surprisingly difficult for existing LLMs. Even when leveraging CoT, they often arrive at the wrong answer. On the other hand, with DiPT, the model would first generate potential methods to solve this question, such as the “arithmetic series method,” “symmetry method,” and “direct summation method” (see details in Figure 3). We provide the comparison of each method’s prompts in Figure 2. While some of these methods may still lead to incorrect answers, others would guide the model to the correct solution. By considering multiple approaches, the model can reflect upon its decisions and ultimately choose the correct answer. This example highlights the potential of DiPT to improve the model’s problem-solving capabilities, even for challenging questions that existing LLMs struggle with.
Figure 2: The comparison of prompts between standard, CoT, and DiPT+ CoT prompting.
Overall, one of the key benefits of DiPT is its potential for error tolerance. In traditional single-path reasoning, if the model’s chosen perspective or solution method suffers from hallucination, it is likely to lead to an incorrect final answer. However, by considering multiple perspectives, DiPT reduces the risk of relying on a single flawed reasoning path. As long as at least one of the considered perspectives or methods leads to the correct solution, the model has the opportunity to self-correct and arrive at the right answer.
3.2 DiPT for Improving Training Data Quality
The key idea behind applying DiPT to improve data quality is to augment the instruction dataset with rationales from multiple solution paths. The intuition behind this approach is that learning from rationales leads to a better mastery of relevant skills and knowledge required to solve a question. When a model is trained on data accompanied by explanations from different perspectives, it can better understand the underlying concepts and principles, rather than simply memorizing associations between inputs and outputs. In contrast, learning directly from raw data may suffer from memorization of associations without proper generalization, leading to poor performance on unseen examples, especially out-of-distribution examples. We verify this in more detail in Table 5, where learning with just instruction-response pairs might sometimes lead to lower performance on other (out-of-distribution) tasks within the same domain, yet learning with rationales always improves out-of-distribution generalization (on average).
To put this idea into practice, we first prompt off-the-shelf models to generate rationales from multiple perspectives for each question in the instruction dataset using the approach detailed in Section 3.1. We then replace the original instruction dataset responses with the corresponding generated responses containing multiple solution paths leading to the answers. We then fine-tune the model on this augmented data.
4 Experiment
This section presents experiments designed to investigate the following questions: (1) How does the integration of perspective-taking into existing reasoning methods impact their performance across various tasks? We evaluate its effect on both the accuracy and robustness to paraphrased problem statements. (Section 4.1) (2) What novel applications can be developed to harness DiPT’s advanced context understanding and accurate reasoning capabilities? Specifically, we will explore its potential in harmful query moderation and dataset error detection. (Section 4.2). (3) How does fine-tuning models on datasets enriched with perspective-taking affect their performance on both in-distribution and out-of-distribution tasks? (Section 4.3).
By addressing these questions, we aim to provide a comprehensive evaluation of the proposed approach and offer insights into its effectiveness, versatility, and generalizability.
4.1 DiPT Integration Impact on Inference
To understand the impact of perspective-taking on reasoning, we demonstrate the effect of adding DiPT to diverse reasoning methods. We considered three existing methods: CoT, which performs step-by-step reasoning; Rephrase and Respond (RaR), which rephrases and expands the question; and Analogical Reasoners (ANL), which self-generates examples similar to the problem. This diverse set allows us to assess the generalizability of DiPT across different reasoning paradigms. We emphasize that the goal is not to exhaustively evaluate DiPT with every state-of-the-art method. Instead, our focus is to understand the specific impact of perspective-taking on reasoning performance.
Experimental Setup.
We perform inference-stage experiments on $7$ tasks: AG News, CosmosQA, RTE, SST-5, SVAMP, TREC, and TruthfulQA. For AG News, SST-5, and TREC, we measure the Top-2 accuracy, as it is possible for an example to belong to multiple classes. For all other tasks, we apply Top-1 accuracy. We refer the reader to Appendix A for further details on datasets. We evaluate performance over $300$ test examples and report the average after $3$ runs. In the main paper, we report results on the GPT-4-Turbo (November) model Achiam et al. (2023), while we provide results on the open-weight model, Mistral7B-Instruct-v0.1 Jiang et al. (2023), in Appendix B.3. Additionally, we provide each DiPT prompt in Appendix A.4. We report 00-shot results of the target model with standard prompting; for each reasoning method, we report results when prompting with and without perspective-taking (DiPT+<Method-Name>) and the difference in the performance ($\Delta$).
CosmosQA TruthfulQA RTE TREC
Standard (0-Shot) $79$ $83$ $87$ $86$
Chain-of-Thought $79$ $85$ $88$ $87$
DiPT+ Chain-of-Thought $82$ $89$ $91$ $93$
[…]
Rephrase and Respond $80$ $83$ $89$ $89$
DiPT+ Rephrase and Respond $83$ $85$ $90$ $94$
[…]
Analogical Reasoning $81$ $84$ $90$ $90$
DiPT+ Analogical Reasoning $84$ $88$ $90$ $94$
[…]
Table 1: Performance comparison between standard prompting, prompting using reasoning method with and without DiPT. Delta performance denotes the performance change when including perspective-taking to reasoning methods.
Result on Accuracy Improvement.
In Table 1, we observe that adding perspective-taking to each of the reasoning methods improves performance in most cases with even $6\%$ increase for CoT in the TREC dataset. We observe performance increases for all cases except for the analogical reasoning with the RTE dataset, where the performance might have reached its peak due to potential labeling errors within the dataset. We will analyze these errors in detail in Section 4.2.
Figure 3: The comparison of responses between Standard, CoT, and DiPT+ CoT prompting.
To gain deeper insights into the positive quantitative results, Figure 3 presents an illustrative example. This example showcases how explicit exploration of multiple solution paths, enabled by DiPT in conjunction with CoT prompting, allows the language model to self-correct. Standard prompting and CoT prompting typically guide the model along a single path, increasing its susceptibility to errors (as seen in Figure 3, where the answers following their corresponding solution paths are incorrect). Conversely, DiPT prompts the model to explore alternative solutions. This capability allows for robust analysis and comparison of answers, ultimately leading the model to identify and correct errors, resulting in a correct final answer (shown in full in Appendix B.4).
Result on Stable Generation.
While current reasoning methods enhance the model’s capabilities, they may generate erroneous reasoning steps across various problem formulations, as noted in studies by Wang et al. (2023b); Lanham et al. (2023); Turpin et al. (2024). We examine whether incorporating perspective-taking into existing methods can enhance stability across different problem paraphrases, thus improving method reliability. To assess this, we evaluate each method’s output stability by measuring its sensitivity to paraphrased prompts. Specifically, we generate five paraphrases of the same queries used in Table 1 and report the mean performance across these iterations. Paraphrasing templates and examples are provided in Appendix A.6. Due to the automatic nature of paraphrasing, a few cases have lost their original meaning due to simplistic rephrasing, resulting in decreased performance across most scenarios in Table 2. However, we observe that all tested methods (CoT, RAR, and ANL) benefit from incorporating perspective-taking. This is evident in two key findings. First, across all methods, incorporating perspective-taking leads to the best overall performance on paraphrased problems. Second, the performance drops for methods with perspective-taking are usually smaller than those without it.
SST-5 CosmosQA RTE
0-Shot 81 $\rightarrow$ 81 (+0) 79 $\rightarrow$ 73 (-6) 87 $\rightarrow$ 83 (-4)
[…]
DiPT+CoT 91 $\rightarrow$ 90 (-1) 82 $\rightarrow$ 80 (-2) 91 $\rightarrow$ 89 (-2)
[…]
DiPT+RAR 90 $\rightarrow$ 89 (-1) 83 $\rightarrow$ 81 (-2) 90 $\rightarrow$ 88 (-2)
[…]
DiPT+ANL 88 $\rightarrow$ 88 (+0) 84 $\rightarrow$ 81 (-3) 90 $\rightarrow$ 88 (-2)
Table 2: Stability results for each method. We rephrase the original prompts to measure the stability of each method. We compare the results with the ones of the original prompts in Table 1.
4.2 Applications of DiPT Integration
Safety Moderation.
The enhanced context understanding achieved by considering multiple viewpoints is beneficial in various application contexts. Here, we demonstrate a specific example of adversarial prompting, where attackers manipulate harmful queries that the model initially rejects, making them appear safe to the model and eliciting inappropriate responses. This issue arises when the model fails to fully comprehend the input context and naively follows the prompt. We demonstrate that perspective-taking enables the model to shift perspectives during output generation, better grasping the user’s intent.
Figure 4: An example of successful prevention of harmful content generation by DiPT moderation.
Figure 4 shows an example where the model successfully prevents harmful content generation, which would otherwise occur. More examples can be found in Appendix B.6. Additionally, Table 3 provides quantitative results comparing our method’s performance against various defense mechanisms, such as paraphrasing, retokenizing (Jain et al., 2023), or summarizing (Zeng et al., 2024). We consider multiple representative attacks, including prompt automatic iterative refinement (PAIR) (Chao et al., 2023), which leverages the LLM to automatically refine the adversarial prompts; greedy coordinate gradient (GCG) (Zou et al., 2023), optimizes prompts with adversarial suffixes to surpass defenses; and persuasive adversarial prompts (PAP) (Zeng et al., 2024), which tries to surpass the model by leveraging persuasive techniques in the prompts.
Our method achieves a $0\%$ attack success rate (ASR) for PAIR and GCG attacks, where ASR calculation is based on keyword matching Zou et al. (2023). While the ASR for PAP is above $0\%$, the generated output might not necessarily be harmful (e.g., a superficial representation of an imaginary weapon for a story). We verify this with a context-aware harmfulness evaluation (Qi et al., 2024) score of 1.44/5 (where 5 is the most harmful/unaligned). This score demonstrates the effectiveness of our moderation. Our method also achieves a similar MT Bench score (Zheng et al., 2023) as the standard model, indicating the successful generation of benign outputs as intended.
PAIR $\downarrow$ GCG $\downarrow$ Persuasion $\downarrow$ MT Bench $\uparrow$
Standard (0-Shot) 92% 92% 92% 8.97
Paraphrase 20 % 0% 60% 7.99
Base Summary 20% 0% 46% 6.51
Tuned Summary 6% 0% 2% 6.65
DiPT 0% 0% 20% 8.97
Table 3: Results of applying defense methods to different attacks by showing the attack success rate (ASR) and the usefulness score (MT Bench) of the model.
Dataset Labeling Error Detection.
We observe that although our method achieves high performance, it cannot reach $100\%$. Upon closer examination, we identified errors within datasets that prevented our method from achieving a perfect score. Current works in reasoning often use datasets commonly employed in NLP. However, for a dataset to serve as a reliable benchmark, it must exhibit high quality without errors. Otherwise, achieving high performance on inaccurately labeled data can mislead the comprehension of the method. Consequently, we aim to prevent such errors in these datasets. We apply our method to verify the labeling of these datasets and identify potential errors in the misalignment of the labels. Specifically, we employ DiPT on the gpt-4-1106-preview model to identify mismatched labels between the predicted and annotated labels. Then, to evaluate the correctness of error identification by our method, we leverage expertise evaluations from several powerful LLMs, including Bard/Gemini and Claude, in conjunction with our judgments. We use fine-grained metrics to better categorize the labeling errors: Wrong label, where all experts disagree with the original ground truth label, Ambiguous examples, where some experts disagree with the original label, and False Positives, where all experts agree with the original label.
SST-5
AG News
TREC
DBPedia
CosmosQA
SVAMP
TruthfulQA
RTE
Wrong 15 4 5 1 5 5 8 3
Ambiguous 4 0 2 0 6 0 3 5
False Positive 1 0 2 0 2 0 10 2
Table 4: Quantitative result of detection of wrong examples found in each of the datasets (over 100 test samples) detected by DiPT.
As shown in Table 4, our method can identify potential incorrect labels, including those ambiguous cases that present challenges for both the model and human assessors. We offer examples of errors in datasets in Appendix B.5. We believe that our method can improve automatic mislabeling detection with enhanced interpretability.
In-Distribution Performance Out-of-Distribution (In-Domain) Performance
Mistral7B
Base
Llama3-8B
Base
Mistral7B
Instruct-v0.2
Lamma3-8B
Instruct
Mistral7B
Base
Lamma3-8B
Mistral7B
Instruct-v0.2
Lamma3-8B
Instruct
OpenbookQA Test Language Understanding and Knowledge
Base Model 43.80 45.00 45.40 43.20 67.68 67.86 67.36 65.99
OpenbookQA
Plain 3K
44.00 44.60 45.40 42.80 67.71 67.82 67.51 65.91
OpenbookQA
CoT 3K
44.20 44.60 45.80 43.00 67.78 68.02 67.70 66.11
OpenbookQA
DiPT+ CoT 3K
44.20 45.00 46.00 43.20 67.84 68.21 68.11 66.21
GSM8K Test Mathematical Reasoning
Base Model 6.60 14.78 21.00 33.74 31.86 47.60 59.67 76.07
GSM8K
Plain 3K
7.73 14.94 21.15 32.65 46.61 35.52 57.99 77.85
GSM8K
CoT 3K
8.91 15.39 25.01 40.38 62.85 58.62 68.81 80.68
GSM8K
DiPT+ CoT 3K
12.96 16.40 24.26 42.50 67.22 69.08 70.64 81.02
CoQA Test Commonsense Reasoning
Base Model 80.68 80.63 76.89 78.13 62.30 68.36 66.99 70.29
CoQA
Plain 3K
80.78 80.75 79.76 78.01 62.60 68.55 66.80 70.13
CoQA
CoT 3K
80.82 80.67 77.98 78.25 62.92 68.96 67.72 70.37
CoQA
DiPT+ CoT 3K
81.19 80.90 79.06 78.35 63.00 69.51 67.87 70.48
XWinograd Test Multilingualism
Base Model 81.46 81.43 79.52 79.68 60.21 62.52 60.74 62.65
XWinograd
Plain 3K
81.43 81.36 81.95 79.46 60.26 62.82 61.53 62.56
XWinograd
CoT 3K
81.50 81.39 81.48 80.00 60.37 63.26 61.63 62.72
XWinograd
DiPT+ CoT 3K
81.50 81.41 82.15 80.06 60.47 63.35 61.90 62.78
Mistral7B
Base
Llama3-8B
Base
Mistral7B
Instruct-v0.2
Lamma3-8B
Instruct
Mistral7B
Base
Lamma3-8B
Mistral7B
Instruct-v0.2
Lamma3-8B
Instruct
Table 5: The fine-tuning results of four different models. The models are trained separately on OpenbookQA, GSM8K, CoQA, and XWinograd and evaluated on their test split (Left: in distribution) and on the associated domain (Right: in domain). Bold means the highest performance, and underlined means the second highest.
4.3 Impact of DiPT-Enriched Fine-Tuning
In addition to enhancing performance during the inference stage, reasoning methods have also been utilized for instruction tuning large language models to improve their ability to follow instructions. Techniques such as chain-of-thought (Kim et al., 2023; Chai et al., 2024) and in-context learning (Longpre et al., 2023) have been successfully incorporated into various datasets for model tuning. In this study, we explore whether data incorporating perspective-taking can be beneficial for model training. Specifically, we concentrate on chain-of-thought data enriched with perspective-taking.
Experimental Setup.
We consider four models for training: Mistral7B-v0.1, Mistral7B-Instruct-v0.2, Llama3-8B, and Llama3-8B-Instruct. These models are fine-tuned on four distinct datasets, each representing a different task domain: OpenbookQA (common knowledge and understanding), GSM8K (grade school math word problems), CoQA (conversational dataset), and XWinograd (multilingual schema challenges). We evaluate the models’ performance in two settings, to assess their in-distribution and out-of-distribution generalization capabilities. For the in-distribution evaluation, we use the respective test split of the training distribution it was trained on to calculate the model’s performance. For the out-of-distribution (in-domain) evaluation, we use other datasets from a similar task domain to evaluate the model’s performance on data outside the training distribution but within the same domain. We group datasets into following domains: language understanding and knowledge (OpenbookQA, MMLU, PIQA, Hellaswag, Lambada), mathematical reasoning (GSM8K, MultiArith, SVAMP, AddSub), commonsense reasoning (CoQA, WSC, Winogrande, ARC-challenge), and multilingualism (XWinograd, WMT16, Lambada-multilingual). We train each model with the original dataset (plain), the CoT version of the dataset, or the DiPT+ CoT version of the dataset, using $3{,}000$ samples for each experiment, except for the base model. For further experimental details and all metrics, please refer to Appendix A.
Results.
We present the results in Table 5. As expected, training the model on the CoT version of the dataset improves performance compared to training on the original dataset, as shown in (Kim et al., 2023). However, our findings reveal that training the model on DiPT+ CoT, which incorporates chain-of-thought reasoning data enhanced with perspective-taking, further enhances performance on downstream tasks across various models. We hypothesize that improving data quality by integrating perspective-taking positively impacts the model’s reasoning capabilities. Interestingly, while direct training on the original dataset might not always yield improvement on out-of-distribution datasets and could even degrade performance, training on rationales, including either CoT or CoT with multi-perspective rationales (DiPT+CoT), consistently improves the average out-of-distribution performance. This observation suggests that rationales might capture shared knowledge across different datasets within the same domain, despite the large variances exhibited by these datasets. Training on a specific dataset might lead to forgetting or overfitting, resulting in poor generalization on other datasets. In contrast, training with rationales could provide a potential pathway to reconcile the conflicts between different datasets, allowing for better generalization and performance across the domain. Therefore, further exploration of perspective-taking in model training is a promising research direction. Additionally, applying DiPT to other reasoning methods might yield similar results, which we leave for future work. We refer to Appendix B for a breakdown of results.
5 Conclusion
In this work, we explore the impact of perspective-taking on reasoning in language models. We investigate whether adding diversified perspective-taking to current reasoning methods can enhance model performance. Our findings show that perspective-taking in generating reasoning improves the model’s understanding of problem context, leading to better answers through corroboration of alternative solutions. Instruction-tuning the model with perspective-taking data further enhances its capabilities compared to chain-of-thought data. We demonstrate the applications of advanced context-understanding capabilities enabled by perspective-taking in the safety and data quality refinement context.
6 Limitations
Despite the improved reasoning capabilities, incorporating diverse perspectives in text generation comes with the cost of extra time. While there are high-stake applications where reasoning accuracy outweighs time costs, there are also scenarios where time constraints might be an important consideration, particularly in real-time applications of LLMs. To address this issue, one potential solution is to adopt an adaptive perspective generation approach. In this approach, the model dynamically adjusts the number of perspectives generated based on the complexity of the problem or the confidence in the initial answer. Another potential fix is to incorporate diverse perspectives during the training phase and then distill the insights gained from multiple perspectives into a more compact model that does not explicitly generate multiple perspectives during inference. However, the effectiveness of these approaches may vary depending on the specific application and the characteristics of the LLM being used. We believe that the in-depth exploration of these ideas is a promising direction for future research.
7 Ethical Considerations
As our method is applied in the model output moderation, it is important to consider the consequences of this mechanism. On one hand, we believe our method can improve the model’s response. However, at the same time, it also controls the generation of harmful responses by the model. It is important to discuss what exactly should be and should not be outputted by the model.
Appendix A Experimental Details & Tasks
A.1 Tasks for Inference Stage
AG News (AG’s News Corpus)
Zhang et al. [2015]. The AG News dataset is a collection of news articles categorically labeled into four classes (World, Sports, Business, and Science/Technology), providing a resource for text classification and topic modeling tasks. As news can belong to more than one category, we use top-2 accuracy.
SST-5 (Stanford Sentiment Treebank)
Socher et al. [2013]. The SST-5 dataset is a sentiment analysis dataset consisting of movie reviews categorized into five sentiment classes, including very negative, negative, neutral, positive, and very positive. We use a top-2 accuracy across methods because a sentiment might lie between 2 neighboring classes due to interpretation.
DBPedia
Auer et al. [2007]. The DBpedia dataset is a knowledge base extracted from Wikipedia, representing structured information about a wide range of entities, including persons, places, organizations, and abstract concepts. We use top-1 accuracy.
CosmosQA (Commonsense Machine Comprehension)
Huang et al. [2019]. The CosmosQA dataset is a reading comprehension dataset requiring contextual commonsense reasoning. The questions are posed as multi-choice problems that ask about likely causes or effects of events. We use top-1 accuracy.
TREC (Text REtrieval Conference)
Li and Roth [2002], Hovy et al. [2001]. The TREC dataset is a question type classification dataset, which contains 6 coarse class labels. We use top-2 accuracy as the question type might belong to more than one category.
SVAMP (Simple Variations on Arithmetic Math word Problems)
Patel et al. [2021a]. The SVAMP dataset is consists of elementary-level math word problems. The dataset consists variations of the problems to test the model’s sensitivity to question understanding. Since the provided dataset is a single-answer dataset, we created three neighboring answers in addition to the groundtruth answer to make the problems multi-choice. We use top-1 accuracy.
TruthfulQA
Lin et al. [2022]. The TruthfulQA dataset is used to measure the truthfulness of the model’s output generation. These problems are prone to be incorrectly answered if fallen into wrong beliefs and require correct pretrained information to be answered. We use top-1 accuracy.
RTE (Recognizing Textual Entailment)
Cooper et al. [1996], Dagan et al. [2005]. The RTE dataset tests the language model in recognizing textual entailment in the provided context. The classification is binary. We use top-1 accuracy.
A.2 Tasks for Fine-Tuning Stage
For these tasks, we use the popular evaluation repository LM Evaluation Harness to evaluate results for the following tasks [Gao et al., 2023].
Language Understanding and Knowledge
  • OpenbookQA [Mihaylov et al., 2018] - dataset designed to evaluate a model’s ability to apply elementary science knowledge to answer questions. We use the normalized top-1 accuracy.
  • MMLU [Mihaylov et al., 2018] - a comprehensive dataset encompassing a wide range of subjects to assess a model’s understanding across various academic disciplines and professional domains. We use the top-1 accuracy.
  • PIQA [Bisk et al., 2020] - a dataset that tests a model’s commonsense knowledge about the physical world. We use the normalized top-1 accuracy.
  • Hellaswag [Zellers et al., 2019] - a challenging dataset for commonsense reasoning, focusing on completing sentences in a way that makes sense in context. We use the normalized top-1 accuracy.
  • LAMBADA [Paperno et al., 2016] - a dataset designed to evaluate the ability of language models to understand and predict a missing word in a passage. We use the top-1 accuracy.
Mathematical Reasoning
  • GSM8K [Cobbe et al., 2021] - a dataset containing $8{,}000$ high-quality grade school math word problems designed to test arithmetic reasoning. We use the normalized 0-shot exact matching (flexible) accuracy.
  • MultiArith [Roy and Roth, 2016] - a dataset focused on arithmetic word problems that require multiple steps to solve. We use the normalized 0-shot exact matching (flexible) accuracy.
  • SVAMP (Single Variable Arithmetic Multiple Problems) [Patel et al., 2021b] - a dataset created to assess the robustness of models on arithmetic word problems. We use the normalized 0-shot exact matching (flexible) accuracy.
  • AddSub [Mishra et al., 2022] - a dataset consisting of arithmetic word problems that involve simple addition and subtraction. We use the normalized 0-shot exact matching (flexible) accuracy.
Commonsense Reasoning
  • CoQA (Conversational Question Answering) [Reddy et al., 2019] - dataset is designed for building conversational question answering systems. We use the F1 score.
  • WSC (The Winograd Schema Challenge)  [Levesque et al., 2012] - a dataset testing commonsense reasoning by identifying pronouns. We use the top-1 accuracy.
  • Winogrande [ai2, 2019] - a dataset extending WSC with more diverse and challenging sentences. We use the top-1 accuracy.
  • ARC Challenge (AI2 Reasoning Challenge) [Clark et al., 2018] - a dataset comprising of difficult multiple-choice science questions. We use the normalized top-1 accuracy.
Multilingualism Reasoning
  • XWinograd [Muennighoff et al., 2022, Tikhonov and Ryabinin, 2021] - a multilingual version of the Winograd Schema Challenge. We use the top-1 accuracy.
  • WMT16 [Bojar et al., 2016] - a dataset consisting of parallel corpora and evaluation data for machine translation tasks. We report CHRF, BLEU, and TER scores and we use the CHRF [Popović, 2015] score accuracy for calculating the domain performance.
  • LAMBADA Multilingual [Paperno et al., 2016] - a dataset extending the original LAMBADA dataset to multiple languages. We use the top-1 accuracy.
A.3 Hyperparameters
In our fine-tuning experiments, we train four models Mistral7B-v0.1, Mistral-7B-Instruct-v0.2 [Jiang et al., 2023], Llama3-8B, and Llama3-8B-Instruct [AI@Meta, 2024] with three Nvidia A100 80G GPUs. We follow the hyperparameter setup from Ethayarajh et al. [2023]. As such, we use a batch size of 32 and train for a single epoch. We keep the learning rate to be $75e-$ 7 as implemented. The maximum sequence length is set to 2048. We use RMSprop as our optimizer with warmup stages for 150 steps. The mixed precision is bfloat16.
A.4 Task Prompts
In this section, we provide the general format of the prompts for each dataset we have implemented:
  • AG News
    AG News Given the news article: {news article} Which two of the following cate- gories the article belongs to: World or Sport or Business or Science/Technology? {method prompt}
  • SST-5
    SST-5 Given the review: {review} Which two of the following sen- timents the review belongs to: very positive or positive or neutral or negative or very negative? {method prompt}
  • DBPedia
    DBPedia Given the subject with a description: subject: {review} description: {description} Which category the subject belongs to: Company or Educational Institution or Artist or Athlete or Office Holder or Mean Of Transportation or Building or Natural Place or or Building or Natural Place or Village or Animal or Plant or Album or Film or Written Work? {method prompt}
  • CosmosQA
    CosmosQA Given a context: {context} Question: {question} Choose the answer from below: 1: {option 1} 2: {option 2} 3: {option 3} 4: {option 4} {method prompt}
  • TREC
    TREC Given the question: {question} Give the category of the question: Abbreviation or Entity or Description and abstract concept or Human being or Location or Numeric value. {method prompt}
  • SVAMP
    SVAMP Given a scenario: {scenario} Question: {question} Choose the answer from below: 1: {option 1} 2: {option 2} 3: {option 3} 4: {option 4} {method prompt}
  • TruthfulQA
    TruthfulQA Given a question: {question} Options: 1: {option 1} 2: {option 2} 3: {option 3} 4: {option 4} {method prompt}
  • RTE
    RTE Given a premise: {question} Hypothesis: {hypothesis} Is the given hypothesis a strict entailment of the premise? Yes or No? {method prompt}
A.5 Method Prompts
In this section, we provide the general format of the prompts for each method we have implemented.
A.5.1 Baseline Method Prompts
  • Automatic/0-Shot Chain-of-Thought
    0-Shot CoT {Task prompt} Let’s think step by step.
  • In-Context Learning
    k-Shot ICL {k demonstrations} {Task prompt}
  • Rehprase and Response (RaR)
    RaR {Task prompt} Rephrase and expand the question, and respond.
  • Analogical Reasoners (ANL)
    RaR {Task prompt} Provide relevant problems as examples. Afterward, proceed to solve the initial problem.
  • DiPT+ Default
    DiPT+ Default {Task prompt} Before choosing the answer, for each option explain if it is possible or not. Choose the selected answers.
    For the sentiment analysis, we replace the word “option” with the word “sentiment”.
A.5.2 DiPT  Prompts in Table 1
  • DiPT+ Rephrase and Respond
    DiPT+ Rephrase and Respond {Task prompt} Before choosing the answer, for each option explain if it is possible or not. Rephrase and expand the question, and respond. Choose the selected answers.
  • DiPT+ Chain-of-Thought
    DiPT+ Chain-of-Thought {Task prompt} Before choosing the answer, for each option explain if it is possible or not. Let’s think step by step. Choose the selected answers.
  • DiPT+ Analogical Reasoner
    DiPT+ Analogical Reasoner {Task prompt} Before choosing the answer, for each option explain if it is possible or not. Provide relevant problems as examples. Afterward, proceed to solve the initial problem. Choose the selected answers.
A.6 Paraphrase prompt in Table 2
For the stability experiment in Section 4.1, we have automatically paraphrased the prompts using the gpt-4-1106-preview model and used the following commands for each dataset we implemented:
  • CosmosQA
    Paraphrase the CosmosQA query Paraphrase the following text preserving the structure (Context and Question) and do not answer the question: Context: {context} Question: {question}
  • RTE
    Paraphrase the RTE query Paraphrase the following text preserving the structure (Premise and Hypothesis) and Premise: {premise} Hypothesis: {hypothesis}
  • SST-5
    Paraphrase the SST-5 query Paraphrase the given review preserving its sentiment: Review: {review}
A.7 DiPT  as a moderation mechanism used in Table 3
Here, we provide the adaptation of our method’s prompt into a moderation mechanism, which goal is to prevent harmful output generation while coordinating benign conversation.
DiPT as Moderation Given a prompt: {review} Before answering this prompt, evaluate if your answer would be potentially harmful and explain why? Would it be dangerous and why? Would it be toxic and why? Or elicit sexual content and why? Explain each case and provide a reasoning. Then depending on the severity of your analysis, consider if you should answer the question. If so, provide the appropriate answer to the question.
Appendix B Additional Results
B.1 Quantitative result for fine-tuning with perspective-taking enriched data.
We present a breakdown of out of distribution (in-domain) results for fine-tuning the model with perspective-taking enriched datasets.
Language Understanding and Knowledge
Model Dataset OpenbookQA MMLU PIQA Hellaswag LAMBADA Average
Mistral7B-v0.1 Base Model 43.80 58.73 82.21 81.08 72.60 67.68
OpenbookQA 3K 44.00 58.68 82.23 81.09 72.57 67.71
OpenbookQA CoT 3K 44.20 58.70 82.26 81.26 72.48 67.78
OpenbookQA DiPT + CoT 3K 44.20 58.70 82.43 81.32 72.57 67.84
Llama3-8B Base Model 45.00 62.05 80.74 79.16 72.33 67.86
OpenbookQA 3K 44.60 62.10 80.83 79.43 72.15 67.82
OpenbookQA CoT 3K 44.60 62.04 80.95 79.72 72.81 68.02
OpenbookQA DiPT + CoT 3K 45.00 62.40 80.96 79.78 72.94 68.21
Mistral7B-Instruct-v0.2 Base Model 45.40 58.77 80.52 83.72 68.40 67.36
OpenbookQA 3K 45.40 58.70 80.52 83.62 69.32 67.51
OpenbookQA CoT 3K 45.80 58.71 80.63 83.61 69.78 67.70
OpenbookQA DiPT + CoT 3K 46.00 58.77 80.68 83.69 71.43 68.11
Llama3-8B-Instruct Base Model 43.20 63.85 78.56 75.81 68.54 65.99
OpenbookQA 3K 42.80 63.88 78.58 75.87 68.41 65.91
OpenbookQA CoT 3K 43.00 63.86 78.62 76.59 68.46 66.11
OpenbookQA DiPT + CoT 3K 43.20 63.89 78.73 76.70 68.54 66.21
Table 6: Break Down of Results for Language Understanding and Knowledge
Mathematical Reasoning
Model Dataset GSM8K MultiArith SVAMP AddSub Average
Mistral7B-v0.1 Base Model 6.60 27.20 36.00 57.65 31.86
GSM8K 3K 7.73 76.00 35.10 67.61 46.61
GSM8K CoT 3K 8.91 90.00 68.40 84.09 62.85
GSM8K DiPT+ CoT 3K 12.96 91.00 80.00 84.94 67.22
Llama3-8B Base Model 14.78 32.20 56.00 87.43 47.60
GSM8K 3K 14.94 35.50 35.10 56.53 35.52
GSM8K CoT 3K 15.39 78.80 68.40 71.86 58.62
GSM8K DiPT+ CoT 3K 16.40 91.10 79.90 89.92 69.08
Mistral7B-Instruct-v0.2 Base Model 21.00 69.00 64.00 84.70 59.67
GSM8K 3K 21.15 66.00 70.10 74.71 57.99
GSM8K CoT 3K 25.01 88.30 76.40 85.51 68.81
GSM8K DiPT+ CoT 3K 24.26 91.60 79.50 87.22 70.64
Llama3-8B-Instruct Base Model 33.74 97.00 82.00 91.53 76.07
GSM8K 3K 32.65 98.00 89.00 91.76 77.85
GSM8K CoT 3K 40.38 98.00 90.00 94.33 80.68
GSM8K DiPT+ CoT 3K 42.50 99.00 89.00 94.60 81.02
Table 7: Break Down of Results for Mathematical Reasoning
Commonsense Reasoning
Model Dataset CoQA WSC Winogrande ARC Challenge Average
Mistral7B-v0.1 Base Model 80.68 40.38 73.95 54.18 62.30
CoQA 3K 80.78 40.38 74.11 55.12 62.60
CoQA CoT 3K 80.82 40.38 74.27 56.23 62.92
CoQA DiPT + CoT 3K 81.19 40.38 74.11 53.24 63.00
Llama3-8B Base Model 80.63 66.35 73.24 54.27 68.36
CoQA 3K 80.75 66.35 72.85 54.86 68.55
CoQA CoT 3K 80.67 67.31 73.01 55.38 68.96
CoQA DiPT + CoT 3K 80.90 68.27 73.48 89.92 69.51
Mistral7B-Instruct-v0.2 Base Model 76.89 61.54 73.56 55.97 66.99
CoQA 3K 79.76 61.54 73.01 55.89 66.80
CoQA CoT 3K 77.98 61.54 75.30 56.06 67.72
CoQA DiPT + CoT 3K 79.06 61.54 74.90 55,97 67.87
Llama3-8B-Instruct Base Model 78.13 74.04 71.98 57.00 70.29
CoQA 3K 78.01 74.04 71.98 56.48 70.13
CoQA CoT 3K 78.25 74.04 72.53 56.66 70.37
CoQA DiPT + CoT 3K 78.35 74.04 72.61 56.91 70.48
Table 8: Break Down of Results for Commonsense Reasoning
Multilingualism
Model Dataset XWinograd WMT16 Lambada Multilingual Average
Mistral7B-v0.1 Base Model 81.46 47.31,24.41,68.90 51.87 60.21
XWinograd 3K 81.43 47.43,24.51,69.80 51.93 60.26
XWinograd CoT 3K 81.50 47.61,24.66,68.06 51.99 60.37
XWinograd DiPT + CoT 3K 81.50 47.98,24.90,67.95 51.93 62.52
Llama3-8B Base Model 81.43 55.37, 30.58, 61.19 50.76 62.82
XWinograd 3K 81.36 56.23,31.11,60.17 50.86 63.26
XWinograd CoT 3K 81.39 57.18,31.84,57.94 51.20 63.35
XWinograd DiPT + CoT 3K 81.41 57.40,31.91,57.57 51.25 68.21
Mistral7B-Instruct-v0.2 Base Model 79.52 54.39, 27.89, 60.15 48.31 60.74
XWinograd 3K 81.95 54.35,27.86,60.17 48.29 61.53
XWinograd CoT 3K 81.48 54.61,28.22,59.98 48.79 61.63
XWinograd DiPT + CoT 3K 82.15 54.65,28.22,59.93 48.91 61.90
Llama3-8B-Instruct Base Model 79.68 59.05,32.02,56.67 49.21 62.65
XWinograd 3K 79.46 59.03,31.99,56.63 49.19 62.56
XWinograd CoT 3K 80.00 59.02,32.01,56.58 49.14 62.72
XWinograd DiPT + CoT 3K 80.06 59.04,32.04,56.53 49.23 62.78
Table 9: Break Down of Results for Multilingualism
B.2 Qualitative result showing the output generation structure when applying our method.
We provide an example of the output generated using DiPT as a prompting framework and visualize the modular structure of the comprehensive reasoning and assessment parts.
Figure 5: An example of the output generated by GPT-4-0613 using DiPT on the SST-5 sample. Both modules in the output are divided and colored accordingly. In the first module, the model evaluates each option possibility. In the second module, the model decides on the final answer given above reasoning.
B.3 Results on Mistral 7B
We demonstrate the results on the open-weight model mistral-7b-instruct-v0.1 available on the HuggingFace library, which is a fine-tuned version on instructions of the mistral-7b-v0.1 model Jiang et al. [2023]. We present results in Table 10.
We note that some methods and models might require further careful tuning of the prompt to suit the model. We made our best efforts to find the fitting prompts and we emphasize that identifying effective prompts for each specific model is an important future direction.
0-Shot K-Shot ICL 0-Shot CoT Rephrasing Analogical Reasoner Ours
AG News 77 79 80 70 60 87
CosmosQA 80 82 83 80 74 86
RTE 67 68 69 53 61 80
SST-5 85 82 86 78 84 89
TREC 57 60 62 64 64 68
[…]
B.4 Full Responses to Example in Figure 3
We provide the full responses to the example presented in Figure 3 for each of the method: Figure 6 for standard prompting, Figure 7 for CoT prompting, and Figure 8 for DiPT+CoT prompting. The responses were transcribed from the ChatGPT model and generated on 2024/06/02.
Figure 6: The full response generated with the Standard prompting for the example in Figure 3.
Figure 7: The full response generated with the Chain-of-Thought prompting for the example in Figure 3.
Figure 8: The full response generated with the DiPT+Chain-of-Thought prompting for the example in Figure 3.
B.5 Examples of Mislabeled/Ambiguous Cases in Datasets
We demonstrate potential errors of each studied dataset identified by DiPT. For each example, we present the query, the original ground truth provided by the dataset, and the new recommended label chosen by DiPT. For a dataset to have an error, it can be that the question is ambiguous, the labels are not correct, no one option is definite, or the ground truth is wrong.
  • AG News
    Figure 9: An example of the error in the AG News dataset detected by DiPT and the new label recommended by DiPT.
  • DBPedia
    Figure 10: An example of the error in the DBPedia dataset detected by DiPT and the new label recommended by DiPT.
  • TREC
    Figure 11: An example of the error in the TREC dataset detected by DiPT and the new label recommended by DiPT.
  • RTE
    Figure 12: An example of the error in the RTE dataset detected by DiPT and the new label recommended by DiPT.
  • SST-5
    Figure 13: An example of the error in the SST-5 dataset detected by DiPT and the new label recommended by DiPT.
  • SVAMP
    Figure 14: An example of the error in the SVAMP dataset detected by DiPT and the new label recommended by DiPT.
  • TruthfulQA
    Figure 15: An example of the error in the TruthfulQA dataset detected by DiPT and the new label recommended by DiPT.
  • CosmosQA
    Figure 16: An example of the error in the CosmosQA dataset detected by DiPT and the new label recommended by DiPT.
B.6 Moderation on Potential Harmful Cases
We present a variety of the latest prompts that can lead the model to potentially generate harmful output to the user, which were found online (https://www.jailbreakchat.com/, https://www.reddit.com/r/ChatGPTJailbreak/) discovered by online users. We demonstrate how our moderation mechanism, DiPT, would address these cases in action. We have tested these prompts on ChatGPT for the more capable model. However, these vulnerabilities may have been patched at the time of reading. We hope the presented attacks can offer insight into how to better moderate the model in the future rather than encourage their malicious intent. Nonetheless, we want to show the universality of our moderation mechanism in handling various potentially harmful prompts that exploit (though the list is not exhaustive):
  • Historical/Education Context
    Figure 17: Example of moderating the attack with the historical/educational context.
  • Hypothetical Scenario
    Figure 18: Example of moderating the attack with the hypothetical scenario.
  • Long Context to Disguise as a Character
    Figure 19: Example of moderating the attack with long context (Jailbreaking DAN).
  • Foreign Language as a Proxy
    Figure 20: Example of moderating the attack with foreign language (Zulu) as a proxy.
  • Unintentional/Unexpected Setting
    While benign users might start a conversation harmlessly, they may (un)intentionally lead it to a potentially harmful path, which might not be detected by the model safeguards. We show a scenario that might occur:
    Figure 21: Example of moderating the unintentionally harmful output.
  • Benign Case
    In addition to potentially harmful cases, we also present a benign scenario, where a user inputs a benign prompt, and we show the behavior of our mechanism toward that prompt:
    Figure 22: Example of moderating a benign input.